Solve the following first order linear equations



1)dy/dx=e^2x+3y



2)xy×dy/dx=2(y+3)



3)(1+x^2).dy/dx-xy(1+y)=0



4)sinx/1+y.dy/dx=cosx



5)cos^2x.dy/dx=y-3




1
Expert's answer
2022-02-10T14:21:15-0500

1)

"\displaystyle
 \frac{dy}{dx}-3y=e^{2x}"

Integrating factor "\displaystyle
=e^{-3x}"

Multiplying both sides of the given DE yields;

"\displaystyle \frac{d}{dx}(ye^{-3x})=e^{-x}\\
ye^{-3x}=\int e^{-x}\ dx=-e^{-x}+c," where c is an arbitrary constant.

"\displaystyle
\Rightarrow y=ce^{3x}-e^{2x}"


2)

"\displaystyle
xy\frac{dy}{dx}=2(y+3)"

By method of separation of variables,

"\displaystyle
\frac{y}{y+3}\frac{dy}{dx}=\frac{2}{x}"

Integrating both sides wrt x yields;

"\displaystyle
\int\frac{y}{y+3}\frac{dy}{dx}\ dx=\int\frac{2}{x}\ dx\\
\Rightarrow \int\frac{y}{y+3}\ dy=\int\frac{2}{x}\ dx\\
\Rightarrow\int\left(1-\frac{3}{y+3}\right)\ dy=\int\frac{2}{x}\ dx\\
\Rightarrow y-3\ln(y+3)=2\ln(x)+\ln(c)", where c is an arbitrary constant

"\displaystyle
\Rightarrow y=cx^2(y+3)^3"


3)

"\displaystyle
(1+x^2)\frac{dy}{dx}-xy(1+y)=0\\
\text{By method of separation of variables,}\\
\Rightarrow \frac{1}{y(y+1)}\frac{dy}{dx}=\frac{x}{1+x^2}\\

\Rightarrow \left(\frac{1}{y}-\frac{1}{y+1}\right)\frac{dy}{dx}=\frac{x}{1+x^2}", by partial fraction

integrating both sides wrt x yields;

"\displaystyle
\int\left(\frac{1}{y}-\frac{1}{y+1}\right)\frac{dy}{dx}\ dx=\int\frac{x}{1+x^2}\ dx\\
\Rightarrow\int\frac{1}{y}\ dy-\int\frac{1}{y+1}\ dy=\int\frac{x}{1+x^2}\ dx\\
\Rightarrow \ln(y)-\ln(y+1)=\frac{1}{2}\ln(x^2+1)+\ln(c)\\", where c is an arbitrary constant

"\displaystyle
\frac{y}{y+1}=c\sqrt{x^2+1}"


4)

"\displaystyle
\frac{\sin x}{1+y}\cdot\frac{dy}{dx}=\cos x\\"

By method of separation of variables,

"\displaystyle
\Rightarrow \frac{1}{1+y}\frac{dy}{dx}=\frac{\cos x}{\sin x}"

integrating both sides wrt x yields;

"\displaystyle
\int\frac{1}{1+y}\frac{dy}{dx}\ dx=\int\frac{\cos x}{\sin x}\ dx\\
\Rightarrow \int\frac{1}{1+y}\ dy=\int\frac{\cos x}{\sin x}\ dx\\
\Rightarrow\ln(y+1)=\ln(\sin x)+\ln(c)", where c is an arbitrary constant

"\displaystyle
\Rightarrow y=c\sin x-1"


5)

"\displaystyle
\cos^2(x)\frac{dy}{dx}=y-3"

By method of separation of variables,

"\displaystyle
\Rightarrow \frac{1}{y-3}\frac{dy}{dx}=\frac{1}{\cos^2 x}"

integrating both sides wrt x yields;

"\displaystyle
\Rightarrow \int\frac{1}{y-3}\frac{dy}{dx}\ dx=\int\frac{1}{\cos^2 x}\ dx\\
\Rightarrow \int\frac{1}{y-3}\ dy=\int\frac{1}{\cos^2 x}\ dx\\
\Rightarrow \int\frac{1}{y-3}\ dy=\int\sec^2 x\ dx\\
\Rightarrow \ln(y-3)=\tan x+c,"where c is an arbitrary constant.

"\displaystyle
\Rightarrow y=3+Ae^{\tan (x)}", where "A=e^c"